# A Numerical Study on Gauge Symmetry of Electroweak Amplitudes

### v2

@@ III Direct Test of Gauge Symmetry
In this section, we directly test the gauge symmetry in EW scattering amplitudes using the MWI (6). We take the EW processes $W^{+}W^{-}\rightarrow t\bar{t}$ and $W^{+}W^{-}\rightarrow W^{+}W^{-}$ as two examples, replacing the polarization vectors of $W$ bosons in the amplitudes with the 5-component momenta $k^{(*)M}$ and computing $k^{M}\mathcal{M}_{M}$ utilizing HELAS in the GE representation.
In Fig. LABEL:fig:MWI-ww2tt, we show $k^{M}\mathcal{M}_{M}$ for $W^{+}W^{-}\rightarrow t\bar{t}$ as function of $\cos\theta$, where $\theta$ is the scattering angle in the center-of-mass (CM) system. The CM energy and the azimuthal angle are fixed as $\sqrt{s}=1\leavevmode\nobreak\ \mathrm{TeV}$ and $\phi=0$. In Figs. LABEL:fig:MWI-ww2tt(a) and LABEL:fig:MWI-ww2tt(b), the polarization vector of $W^{+}$ is replaced by $k^{M}$, with the helicities of $t\bar{t}$ fixed as ++-+ - and the helicity of $W^{-}$ set to be 0 (longitudinal polarization) and $+$, respectively. In Figs. LABEL:fig:MWI-ww2tt(c) and LABEL:fig:MWI-ww2tt(d), the setup is similar, expect the helicities of $t\bar{t}$ are fixed as +++++ +. The blue solid lines represent the results including all tree-level diagrams, exactly demonstrating the MWI $k^{M}\mathcal{M}_{M}=0$. The red dotted and green dashed lines correspond to the results excluding the $t$-channel diagram and those solely involving the $t$-channel diagram, respectively. These contributions are precisely opposite to each other, indicating a large and exact cancellation between different diagrams as expected.
In Fig. LABEL:fig:MWI-ww2ww, we present the results for $W^{+}W^{-}\rightarrow W^{+}W^{-}$, where one, two, and four polarization vectors of $W$ bosons are replaced by one, two, and four 5-component momenta, respectively. The helicities of the remaining $W$ bosons are uniformly set to $+$. The blue lines represent the sums of all tree-level diagrams, while the red dotted, green dashed, and black dot-dashed lines correspond to the results for the contact, $s$-channel, and $t$-channel diagrams, respectively. Once again, we observe that the MWI is satisfied when all tree-level diagrams are included, confirming the gauge symmetry.
@@ IV Testing Gauge Symmetry through Anomalous Couplings
In the previous section, we confirmed that the gauge symmetry represented by the MWI ensures precise cancellation among various diagrams when one or multiple polarization vectors $\epsilon^{M}(k)$ are replaced by 5-component momenta $k^{M}$. In this section, we take a different approach by modifying certain couplings to examine how the MWI is affected. Such a test of gauge symmetry through anomalous couplings can provide deeper insights into how different couplings and parameters are interrelated through gauge symmetry.
In the GE representation, any vertex involving massive vector bosons contains multiple couplings. This allows us not only to modify the couplings of individual vertices as a whole but also to adjust specific parameters within those couplings. The latter will be studied in Subsection IV.1 with 3-point vertices, while the former will be discussed in Subsection IV.2 with 4-point vertices.
IV.1 3-point Vertices
In this subsection, we examine gauge symmetry through the analysis of 3-point vertices, specifically focusing on the $VVh$, $ff^{\prime}V$, and $VVV$ vertices within the GE representation.
$VVh$ vertex
@@
The processes we choose to test gauge symmetry with the $WWh$ vertex are $W^{+}W^{-}\rightarrow t\bar{t}$ and $W^{+}W^{-}\rightarrow hh$. We first compare the cross sections vs. $\delta_{WWh}^{23,4}$ in both the unitary gauge and the GE representation, the results are shown in Fig. LABEL:fig:wwh_gauge. Just as expected, when $\delta^{23}_{WWh}$ and $\delta^{4}_{WWh}$ are nonzero, the cross sections in the GE representation deviate from those in the unitary gauge, which remain unchanged and retain the SM values. The larger $\delta^{23}_{WWh}$ and $\delta^{4}_{WWh}$ are, the larger are the deviations. Compared to $WW\rightarrow hh$, $WW\rightarrow t\bar{t}$ exhibits significantly lower sensitivity to $\delta^{23}_{WWh}$ and $\delta^{4}_{WWh}$. We have to plot $\delta^{23}_{WWh}$ and $\delta^{4}_{WWh}$ over the range [100,100][-100,100][ - 100 , 100 ] for $WW\rightarrow t\bar{t}$, but only [1,1][-1,1][ - 1 , 1 ] for $WW\rightarrow hh$. It is not surprising, given that the $WWh$ vertex contributes to three channels of $WW\rightarrow hh$, but only one in $WW\rightarrow t\bar{t}$. Another notable observation is that the sensitivity of cross section is much higher to $\delta^{23}_{WWh}$ than $\delta^{4}_{WWh}$ for $WW\rightarrow hh$, but the sensitivities are approximately the same for $WW\rightarrow t\bar{t}$.
Next, we compute $k_{M}\mathcal{M}^{M}$ or $k_{1M}k_{2N}\mathcal{M}^{MN}$ for these processes as a way of testing gauge symmetry by replacing polarization vectors of one or two $W$ bosons with 5-component momenta. The results are shown in Fig. LABEL:fig:WI-ww2hh for $WW\rightarrow hh$ and Fig. LABEL:fig:WI-ww2tt for $WW\rightarrow t\bar{t}$, respectively. For $WW\rightarrow t\bar{t}$, we analyze various helicity combinations. In certain helicity configurations, the results consistently vanish due to the violation of angular momentum conservation. Apart from these cases, the anomalous couplings lead to violations of the MWI. For $WW\rightarrow t\bar{t}$, the deviations for $\delta^{23}_{WWh}$ are larger than those for $\delta^{4}_{WWh}$.
@@
The $ff^{\prime}V$ vertex has two sub-vertices: $ff^{\prime}V$ and $ff^{\prime}\varphi$. The $ff^{\prime}V$ coupling can be parameterized as $-i\gamma^{\mu}(g_{\mathrm{L}}P_{\mathrm{L}}+g_{\mathrm{R}}P_{\mathrm{R}})$, while the $ff^{\prime}\varphi$ coupling is $y_{\mathrm{L}}P_{\mathrm{L}}+y_{\mathrm{R}}P_{\mathrm{R}}$. Just like the $VVh$ case, gauge couplings and Yukawa couplings are not independent, as the latter can be expressed in terms of the former.
For a charged current vertex such as $u\bar{d}W$, we have
@@
For a neutral current vertex such as $u\bar{u}Z$, we have
@@
with $s_{\mathrm{W}}\equiv\sin\theta_{\mathrm{W}}$ and $c_{\mathrm{W}}\equiv\cos\theta_{\mathrm{W}}$. To sum up, the $ff^{\prime}\varphi$ couplings are essentially Yukawa couplings $y_{f}$, which are related to the fermion masses $m_{f}$ by $y_{f}=gm_{f}/(\sqrt{2}m_{W})$. This relation is protected by gauge symmetry, and therefore the breaking of gauge symmetry will be reflected on violation of this relation.
Similar to $VVh$, we test gauge symmetry by modifying the $ff^{\prime}V$ couplings as following:
@@
The process we choose to examine gauge symmetry for the $ff^{\prime}V$ vertex is $WW\rightarrow t\bar{t}$. This process is ideal because its $t$-channel diagram includes the $t\bar{b}W$ vertex, while its $s$-channel diagram includes the $t\bar{t}Z$ vertex, allowing us to examine both charged and neutral currents.
In Fig. LABEL:fig:ffV_ww2tt_gauge, we demonstrate how the $WW\rightarrow t\bar{t}$ cross sections are changed when the couplings $y_{ttZ}^{\mathrm{L}/\mathrm{R}}$ and $y_{tbW}^{b/t}$ are modified in both the unitary gauge and the GE representation. Similarly, we show how $k_{M}\mathcal{M}^{M}$ or $k_{1M}k_{2N}\mathcal{M}^{MN}$ change with the anomalous $tbW$ couplings in Fig. LABEL:fig:tbW_ww2tt_anom, and with the anomalous $ttZ$ coupling in Fig. LABEL:fig:ttZ_ww2tt_anom, respectively. As the figures show, gauge symmetry is broken when one of the components of $tbW/ttZ$ couplings are modified, both in terms of cross sections and the MWI. For $ttZ$, the sensitivity of $k_{M}\mathcal{M}^{M}$ and $k_{1M}k_{2N}\mathcal{M}^{MN}$ to $\delta_{ttZ}^{\mathrm{L}}$ is much higher than $\delta_{ttZ}^{\mathrm{R}}$. This is expected, since the SM $ttZ$ vertex is dominated by the left-handed interaction. For $tbW$, the sensitivity to $\delta_{tbW}^{t}$ is much higher than $\delta_{tbW}^{b}$. The reason is also not hard to see: the top Yukawa coupling is much larger than the bottom Yukawa coupling.
$VVV$ vertex
In the SM, there two $VVV$ vertices, $W^{+}W^{+}Z$ and $W^{+}W^{-}A$, either of which has 3 types of sub-vertices: $VVV$, $V\varphi V$, and $\varphi\varphi V$. There is no $\varphi\varphi\varphi$-type vertex.
@@
with $c_{2\mathrm{W}}\equiv\cos 2\theta_{\mathrm{W}}$ and $e=gs_{\mathrm{W}}$. Following the analysis for the $VVh$ and $ff^{\prime}V$ vertices, we modify the couplings to test gauge symmetry as following:
@@
We examine gauge symmetry of the $WWZ$ vertex with both the $W^{+}W^{+}\rightarrow W^{+}W^{-}$ and $WW\rightarrow t\bar{t}$ processes. In Figs. LABEL:fig:wwz_ww2ww_anom and LABEL:fig:wwz_ww2tt_anom, we test the MWI by modifying $\delta_{WWZ}^{234}$ and $\delta_{WWZ}^{567}$ for $W^{+}W^{+}\rightarrow W^{+}W^{-}$ and $WW\rightarrow t\bar{t}$, respectively. In both cases, the sensitivity of the violation of the MWI to $\delta_{WWZ}^{567}$ is higher than $\delta_{WWZ}^{234}$, sometimes significantly so. This is somewhat surprising, since the $VV\varphi$-type vertex that $\delta_{WWZ}^{567}$ modifies is suppressed by $e=gs_{\mathrm{W}}$. However, they are typically proportional to $m_{W}$, providing an enhancement that effectively counterbalances the suppressive influence of the weak mixing angle $\theta_{\mathrm{W}}$.
IV.2 4-point Vertices
@@
Gauge symmetry is not only reflected in the couplings within a single vertex, but also in the relations among different vertices. For example, in the $WW\rightarrow WW$ process, where vertices such as $WWWW$, $WWZ$, $WW\gamma$, and $WWh$ are involved, the MWI requires precise relations among their couplings. Any deviation from the SM values breaks gauge symmetry, similar to how it violate unitarity in the gauge representation. We will modify the overall couplings of the vertices to test the MWI. The processes we choose are $W^{+}W^{-}\rightarrow W^{+}W^{-}$, $W^{+}W^{-}\rightarrow hh$, and $W^{+}W^{-}\rightarrow t\bar{t}$.
@@
For $W^{+}W^{-}\rightarrow W^{+}W^{-}$, we separately vary the overall couplings of the $WWh$, $WWZ$, and $WWWW$ vertices and compute $k_{M}\mathcal{M}^{M}$. In Fig. LABEL:fig:4pt_ww2ww, we show the results, listing different helicity combinations in the process. As expected, we can see the violation of the MWI when the anomalous couplings are nonzero. This violation is most sensitive to $\delta_{WWZ}$ and $\delta_{WWWW}$.
@@
One interesting observation is that, in most helicity combinations, $k_{M}\mathcal{M}^{M}$ remains zero or close to zero. This reminds us to be careful in selecting helicities when testing gauge symmetry of massive amplitudes. The exact vanishing of $k_{M}\mathcal{M}^{M}$ in the presence of an anomalous $\delta_{WWh}$ appears to be due to angular momentum conservation: the amplitude vanishes when angular momentum conservation is violated, so only helicity combinations that conserve angular momentum can yield nonzero values.
@@
For $WW\rightarrow hh$, we separately modify the couplings of the $hhh$, $WWh$, and $WWhh$ vertices, and compute $k_{M}\mathcal{M}^{M}$. The results are shown in Fig. LABEL:fig:4pt_ww2hh.
@@
For $WW\rightarrow t\bar{t}$, we separately modify the couplings of the $tbW$, $ttZ$, and $tth$ vertices. The results are shown in Fig. LABEL:fig:4pt_ww2tt.
@@
The overall pattern is similar for all three processes. Angular conservation forbids some helicity combinations, giving $k_{M}\mathcal{M}^{M}=0$ automatically, no matter how the couplings of individual vertices are modified. This, however, does not mean that the MWI is valid regardless of the couplings. Because there are always some helicity combinations, in which a modification of one coupling results in the violation of the MWI.

### v3

@@ III Direct Test of Gauge Symmetry
In this section, we directly test the gauge symmetry in EW scattering amplitudes using the MWI (6). We consider the EW processes $W^{+}W^{-}\rightarrow t\bar{t}$ and $W^{+}W^{-}\rightarrow W^{+}W^{-}$ as two examples, replacing the polarization vectors of $W$ bosons in the amplitudes with the five-component momenta $k^{(*)M}$ and computing $k^{M}\mathcal{M}_{M}$ utilizing HELAS in the GE representation.
In Fig. 6(a), we show $k^{M}\mathcal{M}_{M}$ for $W^{+}W^{-}\rightarrow t\bar{t}$ as a function of $\cos\theta$, where $\theta$ is the scattering angle in the center-of-mass (CM) system. The CM energy and azimuthal angle are fixed as $\sqrt{s}=1\penalty 10000\ \mathrm{TeV}$ and $\phi=0$, respectively. In Figs. 6(a)(a) and 6(a)(b), the polarization vector of $W^{+}$ is replaced by $k^{M}$, with the helicities of $t\bar{t}$ fixed as ++- and the helicity of $W^{-}$ set to be 0 (longitudinal polarization) and ++. In Figs. 6(a)(c) and 6(a)(d), the setup is similar, except the helicities of $t\bar{t}$ are fixed as ++++. The blue solid lines represent the results including all tree-level diagrams, demonstrating the MWI $k^{M}\mathcal{M}_{M}=0$. The red dotted and green dashed lines correspond to the results excluding the $t$-channel diagram and those solely involving the $t$-channel diagram, respectively. These contributions are precisely opposite to each other, indicating a large and exact cancellation between different diagrams, as expected.
(a) Testing the MWI by computing $k^{M}\mathcal{M}_{M}$ for $W^{+}W^{-}\rightarrow t\bar{t}$.
In Fig. 6(b), we present the results for $W^{+}W^{-}\rightarrow W^{+}W^{-}$, where one, two, and four polarization vectors of $W$ bosons are replaced by one, two, and four five-component momenta, respectively. The helicities of the remaining $W$ bosons are uniformly set to ++. The blue lines represent the sums of all tree-level diagrams, whereas the red dotted, green dashed, and black dot-dashed lines correspond to the results for the contact, $s$-channel, and $t$-channel diagrams, respectively. Once again, we observe that the MWI is satisfied when all tree-level diagrams are included, confirming the gauge symmetry.
(b) Testing the MWI by computing $k_{M}\mathcal{M}^{M}$ (a), $k_{1M}k_{2N}\mathcal{M}^{MN}$ (b), and $k_{1M}k_{2N}k^{*}_{3O}k^{*}_{4P}\mathcal{M}^{MNOP}$ for $W^{+}W^{-}\rightarrow W^{+}W^{-}$.
@@ IV Testing Gauge Symmetry through Anomalous Couplings
In the previous section, we confirmed that the gauge symmetry represented by the MWI ensures precise cancellation among various diagrams when one or multiple polarization vectors $\epsilon^{M}(k)$ are replaced by five-component momenta $k^{M}$. In this section, we take a different approach by modifying certain couplings to examine how the MWI is affected. Such a test of gauge symmetry through anomalous couplings can provide deeper insights into how different couplings and parameters are interrelated through gauge symmetry.
In the GE representation, any vertex involving massive vector bosons contains multiple couplings. This allows us not only to modify the couplings of individual vertices as a whole but also to adjust specific parameters within those couplings. The latter will be studied in Subsection IV.1 with three-point vertices, whereas the former will be discussed in Subsection IV.2 with four-point vertices.
IV.1 Three-point Vertices
In this subsection, we examine gauge symmetry through the analysis of three-point vertices, specifically focusing on the $VVh$, $ff^{\prime}V$, and $VVV$ vertices within the GE representation.
VVhVVh vertex
@@ IV.1 Three-point Vertices
(c) Cross sections of $WW\rightarrow hh$ (upper panels) and $WW\rightarrow t\bar{t}$ (lower panels) at $\sqrt{s}=10\penalty 10000\ \mathrm{TeV}$ with varying $\delta^{23}_{WWh}$ and $\delta^{4}_{WWh}$ in the unitary gauge (blue solid lines) and GE representation (red dotted lines).
The processes we choose to test gauge symmetry with the $WWh$ vertex are $W^{+}W^{-}\rightarrow t\bar{t}$ and $W^{+}W^{-}\rightarrow hh$. We first compare the cross sections with respect to $\delta_{WWh}^{23,4}$ in both the unitary gauge and the GE representation, the results are shown in Fig. 6(c). As expected, when $\delta^{23}_{WWh}$ and $\delta^{4}_{WWh}$ are nonzero, the cross sections in the GE representation deviate from those in the unitary gauge, which remain unchanged and retain the SM values. The larger the values of $\delta^{23}_{WWh}$ and $\delta^{4}_{WWh}$, the larger the deviations. Compared with $W\rightarrow hh, WW$$W\rightarrow t\bar{t}$ exhibits significantly lower sensitivity to $\delta^{23}_{WWh}$ and $\delta^{4}_{WWh}$. We must plot $\delta^{23}_{WWh}$ and $\delta^{4}_{WWh}$ over the range [100,100][-100,100] for $WW\rightarrow t\bar{t}$ but only [1,1][-1,1] for $WW\rightarrow hh$. This is not surprising, given that the $WWh$ vertex contributes to three channels of $WW\rightarrow hh$ but only one in $WW\rightarrow t\bar{t}$. Another notable observation is that the sensitivity of the cross section is much higher to $\delta^{23}_{WWh}$ than to $\delta^{4}_{WWh}$ for $WW\rightarrow hh$, but the sensitivities are approximately the same for $WW\rightarrow t\bar{t}$.
(d) $k_{M}\mathcal{M}^{M}$ and $k_{1M}k_{2N}\mathcal{M}^{MN}$ as functions of $\delta^{23}_{WWh}$ and $\delta^{4}_{WWh}$ for $WW\rightarrow hh$ with one or two $W$ polarization vectors are replaced by five-component momenta.
Next, we compute $k_{M}\mathcal{M}^{M}$ or $k_{1M}k_{2N}\mathcal{M}^{MN}$ for these processes as a way of testing gauge symmetry by replacing the polarization vectors of one or two $W$ bosons with five-component momenta. The results are shown in Fig. 6(d) for $WW\rightarrow hh$ and Fig. 6(e) for $WW\rightarrow t\bar{t}$. For $WW\rightarrow t\bar{t}$, we analyze various helicity combinations. In certain helicity configurations, the results consistently vanish due to the violation of angular momentum conservation. Apart from these cases, the anomalous couplings lead to violations of the MWI. For $WW\rightarrow t\bar{t}$, the deviations for $\delta^{23}_{WWh}$ are larger than those for $\delta^{4}_{WWh}$.
(e) $k_{M}\mathcal{M}^{M}$ and $k_{1M}k_{2N}\mathcal{M}^{MN}$ as functions of $\delta^{23}_{WWh}$ and $\delta^{4}_{WWh}$ for $WW\rightarrow t\bar{t}$ with one or two $W$ polarization vectors replaced by five-component momenta. The results for various helicity combinations are shown.
@@
The $ff^{\prime}V$ vertex has two sub-vertices: $ff^{\prime}V$ and $ff^{\prime}\varphi$. The $ff^{\prime}V$ coupling can be parameterized as $i\gamma^{\mu}(g_{\mathrm{L}}P_{\mathrm{L}}+g_{\mathrm{R}}P_{\mathrm{R}})$, whereas the $ff^{\prime}\varphi$ coupling is $y_{\mathrm{L}}P_{\mathrm{L}}+y_{\mathrm{R}}P_{\mathrm{R}}$. Similar to the $VVh$ case, gauge couplings and Yukawa couplings are not independent, as the latter can be expressed in terms of the former.
For a charged current vertex, such as $u\bar{d}W$, we have
@@
For a neutral current vertex, such as $u\bar{u}Z$, we have
@@
with $s_{\mathrm{W}}\equiv\sin\theta_{\mathrm{W}}$ and $c_{\mathrm{W}}\equiv\cos\theta_{\mathrm{W}}$. To sum up, the $ff^{\prime}\varphi$ couplings are Yukawa couplings $y_{f}$, which are related to the fermion masses $m_{f}$ by $y_{f}=gm_{f}/(\sqrt{2}m_{W})$. This relation is protected by gauge symmetry, and therefore, the breaking of gauge symmetry will be reflected in the violation of this relation.
Similar to $VVh$, we test gauge symmetry by modifying the $ff^{\prime}V$ couplings as follows:
@@
The process we choose to examine gauge symmetry for the $ff^{\prime}V$ vertex is $WW\rightarrow t\bar{t}$. This process is ideal because its $t$-channel diagram includes the $t\bar{b}W$ vertex, whereas its $s$-channel diagram includes the $t\bar{t}Z$ vertex, allowing us to examine both charged and neutral currents.
(f) Cross sections of $WW\rightarrow t\bar{t}$ in the unitary gauge and GE representation as functions of the anomalous couplings of ttZttZ (upper panels) and tbWtbW (lower panels).
In Fig. 6(f), we demonstrate how the $WW\rightarrow t\bar{t}$ cross sections are changed when the couplings $y_{ttZ}^{\mathrm{L}/\mathrm{R}}$ and $y_{tbW}^{b/t}$ are modified in both the unitary gauge and GE representation. Similarly, we show how $k_{M}\mathcal{M}^{M}$ or $k_{1M}k_{2N}\mathcal{M}^{MN}$ changes with the anomalous $tbW$ couplings in Fig. 6(g), and with the anomalous $ttZ$ coupling in Fig. 6(h). As the figures show, gauge symmetry is broken when one of the components of $tbW/ttZ$ couplings is modified, both in terms of cross sections and the MWI. For $ttZ$, the sensitivity of $k_{M}\mathcal{M}^{M}$ and $k_{1M}k_{2N}\mathcal{M}^{MN}$ to $\delta_{ttZ}^{\mathrm{L}}$ is much higher than that to $\delta_{ttZ}^{\mathrm{R}}$. This is expected, as the SM $ttZ$ vertex is dominated by the left-handed interaction. For $tbW$, the sensitivity to $\delta_{tbW}^{t}$ is much higher than that to $\delta_{tbW}^{b}$. The reason is also not difficult to understand: the top Yukawa coupling is much larger than the bottom Yukawa coupling.
(g) Testing the MWI by computing $k_{M}\mathcal{M}^{M}$ (upper panels) and $k_{1M}k_{2N}\mathcal{M}^{MN}$ (lower panels) with anomalous $tbW$ couplings for the process $WW\rightarrow t\bar{t}$. Various helicity combinations are shown.
(h) Testing the MWI by computing $k_{M}\mathcal{M}^{M}$ (upper panels) and $k_{1M}k_{2N}\mathcal{M}^{MN}$ (lower panels) with anomalous $ttZ$ couplings for the process $WW\rightarrow t\bar{t}$. Various helicity combinations are shown.
VVVVVV vertex
In the SM, there are two $VVV$ vertices, $W^{+}W^{+}Z$ and $W^{+}W^{-}A$, either of which has three types of sub-vertices: $VVV$, $V\varphi V$, and $\varphi\varphi V$. There is no $\varphi\varphi\varphi$-type vertex.
@@
with $c_{2\mathrm{W}}\equiv\cos 2\theta_{\mathrm{W}}$ and $e=gs_{\mathrm{W}}$. Following the analysis for the $VVh$ and $ff^{\prime}V$ vertices, we modify the couplings to test gauge symmetry as follows:
@@
(i) Testing the MWI by computing $k_{1M}k_{2N}\mathcal{M}^{MN}$ (upper panels), and $k_{1M}k_{2N}k^{*}_{3O}k^{*}_{4P}\mathcal{M}^{MNOP}$ (lower panels) with anomalous $WWZ$ couplings for the process $W^{+}W^{+}\rightarrow W^{+}W^{-}$. In the upper panels, various helicity combinations are shown.
@@
We examine the gauge symmetry of the $WWZ$ vertex with both the $W^{+}W^{+}\rightarrow W^{+}W^{-}$ and $WW\rightarrow t\bar{t}$ processes. In Figs. 6(i) and 6(j), we test the MWI by modifying $\delta_{WWZ}^{234}$ and $\delta_{WWZ}^{567}$ for $W^{+}W^{+}\rightarrow W^{+}W^{-}$ and $WW\rightarrow t\bar{t}$, respectively. In both cases, the sensitivity of the violation of the MWI to $\delta_{WWZ}^{567}$ is higher than that to $\delta_{WWZ}^{234}$, sometimes significantly so. This is somewhat surprising, as the $VV\varphi$-type vertex that $\delta_{WWZ}^{567}$ modifies is suppressed by $e=gs_{\mathrm{W}}$. However, they are typically proportional to $m_{W}$, providing an enhancement that effectively counterbalances the suppressive influence of the weak mixing angle $\theta_{\mathrm{W}}$.
(j) Testing the MWI by computing $k_{M}\mathcal{M}^{M}$ (upper panels) and $k_{1M}k_{2N}\mathcal{M}^{MN}$ (lower panels) with anomalous $WWZ$ couplings for the the process $WW\rightarrow t\bar{t}$. Various helicity combinations are shown.
IV.2 Four-point Vertices
@@
Gauge symmetry is reflected not only in the couplings within a single vertex but also in the relations among different vertices. For example, in the $WW\rightarrow WW$ process, where vertices, such as $WWWW$, WWZWWZ, $WW\gamma$, and $WWh$, are involved, the MWI requires precise relations among their couplings. Any deviation from the SM values breaks gauge symmetry, similar to how it violates unitarity in the gauge representation. We modify the overall couplings of the vertices to test the MWI. The processes we choose are $^{+}W^{-}\rightarrow W^{+}W^{-}, W+W$$^{+}W^{-}\rightarrow hh$, and $W^{+}W^{-}\rightarrow t\bar{t}$.
@@
For $W^{+}W^{-}\rightarrow W^{+}W^{-}$, we separately vary the overall couplings of the $WWh$, WWZWWZ, and $WWWW$ vertices and compute $k_{M}\mathcal{M}^{M}$. In Fig. 6(k), we show the results, listing different helicity combinations in the process. As expected, we can observe the violation of the MWI when the anomalous couplings are nonzero. This violation is most sensitive to $\delta_{WWZ}$ and $\delta_{WWWW}$.
(k) $k_{M}\mathcal{M}^{M}$ for $W^{+}W^{-}\rightarrow W^{+}W^{-}$ as a function of anomalous overall couplings $\delta_{WWh}$ (a), $\delta_{WWZ}$ (b), and $\delta_{WWWW}$ (c). Various helicity combinations are shown.
@@
One interesting observation is that, in most helicity combinations, $k_{M}\mathcal{M}^{M}$ remains zero or close to zero. This reminds us to be careful in selecting helicities when testing the gauge symmetry of massive amplitudes. The exact vanishing of $k_{M}\mathcal{M}^{M}$ in the presence of an anomalous $\delta_{WWh}$ appears to be due to angular momentum conservation: the amplitude vanishes when angular momentum conservation is violated; hence, only those helicity combinations that conserve angular momentum can yield nonzero values.
@@
For $WW\rightarrow hh$, we separately modify the couplings of the hhhhhh, WWhWWh, and $WWhh$ vertices, and compute $k_{M}\mathcal{M}^{M}$. The results are shown in Fig. 6(l).
(l) $k_{M}\mathcal{M}^{M}$ for $WW\rightarrow hh$ as a function of anomalous overall couplings $\delta_{hhh}$ (a), $\delta_{WWh}$ (b), and $\delta_{WWhh}$ (c). Various helicity configurations are shown.
@@
For $WW\rightarrow t\bar{t}$, we separately modify the couplings of the $tbW$, ttZttZ, and tthtth vertices. The results are shown in Fig. 6(m).
(m) $k_{M}\mathcal{M}^{M}$ for $WW\rightarrow t\bar{t}$ as a function of anomalous overall couplings $\delta_{tbW}$ (a), $\delta_{ttZ}$ (b), and $\delta_{tth}$ (c). Various helicity combinations are shown.
@@
The overall pattern is similar for all three processes. Angular conservation forbids some helicity combinations, giving $k_{M}\mathcal{M}^{M}=0$ automatically, regardless of how the couplings of individual vertices are modified. However, this does not imply that the MWI is valid regardless of the couplings, as there are always some helicity combinations in which a modification of one coupling results in the violation of the MWI.